OpenStudy (anonymous):

A 25 ft ladder is leaning against a vertical wall. The bottom of the ladder is pulled horizontally away from the wall at 3 ft/sec. Determine how fast the top of the ladder is sliding when the bottom of the ladder is 15 ft from the wall. Please explain the process!

OpenStudy (anonymous):

 differentiate the function x^2+y^2=25^2 with respect to time 2x dx/dt=-2y dy/dt when x=15 y=20 given dx/dt = 3 find dy/dt which is the speed at which top falls , the negative sign just shows the direction of motion